OpenStudy (safari321):

In 2000, there were about 210,000,000 Americans eligible to vote for President. About 50% of eligible voters did not vote that year. Al Gore got 48% of the votes that were cast. About how many votes did Gore receive?

OpenStudy (anonymous):

(total votes - (totalvotes*0.50) )*0.48 = votes that gore received

OpenStudy (anonymous):

which is the product of "lilg's" answer and 48 percent

OpenStudy (lilg132):

oops didnt even read the bit where it says 48%

OpenStudy (safari321):

so its 48 %

OpenStudy (anonymous):

now you have a formula for next time

OpenStudy (lilg132):

ok so you half the number of voters 210,000,000 / 2 = 105,000,000 people did not vote 105,000,000 / 100 = 1,050,000 (is 1 percent) 1,050,000 * 48 = 50,400,000